OpenStudy (anonymous):

identify the vertex for the graph of y=4(x+3)^2+4

myininaya (myininaya):

y=a(x-h)^2+k vertex is (h,k)

OpenStudy (anonymous):

expand 4(x+3)^2+4 = 4(x^2+6x+9)+4=4x^2+24x+36+4=4x^2+24x+40 vertex = -b/2a=-24/8=-3 plugging back in yielding: (-3,4)

myininaya (myininaya):

wow thats super long anhhuyalex lol

myininaya (myininaya):

but it still works